OpenStudy (anonymous):

what is 28 times 42 ten thousandths? I need help. I can multiply, but I do not understand what 42 ten thousandths means.

OpenStudy (anonymous):

please help. no matter how dumb you think I am I have been going crazy with this one.

OpenStudy (anonymous):

\[28*\frac{42}{10000}=\frac{147}{1250}=0.1176 \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

you bet thank you for helping me.

OpenStudy (anonymous):

\[28*\frac{42}{10000}=\frac{28*42}{10000}=\frac{1176}{10000}=\frac{.1176}{1.0000}=.1176 \]Move the decimal point 4 places to the left so that the denominator is 1.